# Word Embedding Dimension Reduction via Weakly-Supervised Feature Selection

### v1

@@ 4.3 Downstream Tasks
Table 3: Comparison of the results of downstream prediction tasks. The last column is the average of the previous columns.
@@
Table 3 shows the results of various methods on the nine prediction tasks. The last column is the average of the previous columns. Our experiments show that our method outperforms the PCA-based method on most prediction tasks. Also, our method achieves higher average scores than the PCA-based method in all the settings, except for a slightly lower performance when reducing the Fasttext word embedding from 300 to 50 dimensions, with a difference of only 0.02. Our method performs better than the EmbedTextNet model, although it performs slightly worse when reduced to 50 dimensions. However, as shown in the next section, our model is much more efficient and takes less time to train than the EmbedTextNet model. Our method retains more helpful information for most settings while reducing dimensions, achieving a much closer prediction performance to the original 300-dimensional word embeddings. It indicates the effectiveness of our method. Also, it proves that our method can generalize well on various prediction downstream tasks in a zero-shot manner. The Glove word embedding we used was pre-trained on a smaller amount of data and might be vulnerable to noise. Post-processing can help mitigate noise and improve the performance. However, subtle linguistic connections can be crucial when working with prediction tasks. Applying PPA to word embeddings pre-trained on relatively large data may decrease performance.
@@
Table 4 shows the results of various methods on the sentence similarity tasks. Our method is consistently better than the existing methods, except when we reduce Glove embedding to 150 dimensions and Word2vec embedding to 50 dimensions, where the existing methods perform better with less than 0.05 difference. Our method significantly outperforms the existing approach in most sentence similarity tasks, with a gap of up to 8.36. The observation of post-processing is the same as that of the aggregated word similarity dataset. The Word2vec embeddings we use are trained on extensive data and may not benefit much from the PPA.
Table 4: Comparison of the results of downstream sentence similarity tasks. The last column is the average of the previous columns.
@@
In conclusion, our method outperforms existing methods in prediction and similarity tasks, demonstrating its superiority and generalizability.
@@ 4.4 Model Efficiency
We compare the model efficiency of the Algo method, the EmbedTextNet method, and our method when reducing the word embedding dimensions from 300 to 150 in Table 5. Different hardware is used for other methods because of their algorithmic nature. All the experiments were conducted on the same server, equipped with two AMD EPYC 7543 32-core CPUs and seven NVIDIA RTX A6000 GPUs. The deep learning-based model is trained using the GPUs, while the Algo model and our methods are trained using only the CPUs. The EmbedTextNet model takes approximately 30 minutes to train on relatively small GloVe word embeddings.
In contrast, the training time for our WordFS-P-wP and WordFS-S-wP models is less than 19 seconds and 13 seconds, respectively. Our methods without post-processing require even less time — the training of the WordFS-P-woP model is completed within one second. For more significant pre-trained word embeddings, it takes hours to train the EmbedTextNet model. However, our WordFS-P-wP and WordFS-S-wP methods only require 1-2 minutes, making them hundreds of times faster. Our WordFS-P-woP model, without post-processing, consistently completes in around one second, making it thousands or ten thousand times faster than the deep learning-based method.
Additionally, our methods are more efficient than the PCA-based method because they employ PPA before and after the PCA, while we only apply it once before feature selection. The PPA consumes a substantial portion of the training time compared to the feature selection module. Our model, without the post-processing, can achieve even less time. The reason is the post-processing and PCA must be done on the entire vocabulary. However, our feature selection method only focuses on a subset of words that appear in the aggregated word similarity dataset, making the procedure very fast. And the module’s processing time will not be significantly affected by the vocabulary size. In conclusion, Our methods take only a fraction of the time compared to deep learning methods and less than half the time compared to existing PCA-based methods. The experimental results demonstrate that our methods are much more efficient than existing methods.

### v2

@@ 4.3 Downstream Tasks
Table 3: Comparison of the results of downstream prediction tasks, where the last column is the average of the previous columns.
@@
Figure 2: Average accuracy comparison for prediction tasks.
Table 3 shows the results of various methods on the nine prediction tasks. The last column is the average of the previous columns. Figure 2 compares the average accuracy among the top three methods. Experiments show that our method surpasses the UMAP method in all the prediction tasks and outperforms the PCA-based method in most tasks. Also, our method achieves higher average scores than the PCA-based method in all the settings, except for a slightly lower performance when reducing the Fasttext word embedding from 300 to 50 dimensions, with a difference of only 0.02. Our method also performs better than the EmbedTextNet model, although it performs slightly worse when reduced to 50 dimensions. However, as shown in the next section, our model is much more efficient and takes less time to train than the EmbedTextNet model. Our method retains more helpful information for most settings while reducing dimensions, achieving a much closer prediction performance to the original 300-dimensional word embeddings. It indicates the effectiveness of our method. Also, it proves that our method can generalize well on various prediction downstream tasks in a zero-shot manner. The Glove word embedding we used was pre-trained on a smaller amount of data and might be vulnerable to noise. Post-processing can help mitigate noise and improve the performance. However, subtle linguistic connections can be crucial when working with prediction tasks. Applying PPA to word embeddings pre-trained on relatively large data may decrease performance.
Figure 3: Accuracy comparison for prediction tasks using our WordFS method.
Reducing the dimension of data results in information loss, often leading to decreased performance as the dimension decreases. Figure 3 compares the impact of word vector dimensions on performance across different tasks. The impact of dimension reduction on task performance can vary depending on the pre-trained word embeddings used, but the overall trend remains consistent. Various tasks may lose crucial information at different reduction levels. For example, the TREC task’s performance significantly drops when reduced to 50D compared to 100D, while the performance of most other tasks decreases gradually when reduced to a lower amount. When using Word2vec and Fasttext word embeddings, the TREC task’s performance drops significantly when the dimension is reduced to 150D. This suggests that important information about the type of question is lost early on. For most other tasks, performance decreases slightly as the dimension decreases, which is expected when dimensions are reduced. In conclusion, dimension reduction models can preserve essential information for improved overall performance, but the performance of specific tasks may vary without further fine-tuning.
@@
Table 4 shows the results of various methods on the sentence similarity tasks. Figure 2 compares the average Spearman’s rank correlation coefficients among different methods. Our method is consistently better than the existing methods, except when we reduce Glove embedding to 150 dimensions and Word2vec embedding to 50 dimensions, where the existing methods perform better with less than 0.05 difference. Our method significantly outperforms the existing approach in most sentence similarity tasks, with a gap of up to 8.36. The observation of post-processing is the same as that of the aggregated word similarity dataset. The Word2vec embeddings we use are trained on extensive data and may not benefit much from the PPA. The performance of the reduced word embeddings is often as good or even better than the original word vectors for Glove and Fasttext word embeddings in sentence similarity tasks. Even for Word2vec word embeddings, the drop in performance is small. This indicates that our method, which is based on word similarity, effectively retains critical dimensions for sentence similarity tasks while reducing the confusing ones.
Table 4: Performance comparison of downstream sentence similarity tasks, where the last column is the average of the previous columns.
@@
Figure 4: Comparison of the average Spearman’s rank correlation coefficients of sentence similarity tasks.
Figure 5: Spearman’s rank correlation coefficients comparison for similarity tasks using our WordFS method.
Figure 5 compares the impact of word vector dimensions on performance across different tasks. As the dimensions decrease, there is generally a slight drop in performance because some detailed information is discarded. Most lines show a similar trend from 150D to 50D. For the Glove and Word2vec embeddings, our approach significantly improves performance when reducing to 150D. This is understandable because these embeddings are trained on a smaller amount of data, and our approach helps the vectors focus on similarity tasks. Word2vec word embeddings are trained on a much larger amount of data, which is sufficient for them to perform well on sentence similarity tasks.
In conclusion, our method generally outperforms existing methods in prediction and similarity tasks, demonstrating its superiority and generalizability.
@@ 4.4 Model Efficiency
We compare the model efficiency of the Algo method, the EmbedTextNet method, and our method when reducing the word embedding dimensions from 300 to 150 in Table 5. For each pre-trained word embedding, all methods start with the same vocabulary of 300D word vectors and output a file with 150D word vectors. Different hardware is used for other methods because of their algorithmic nature. All the experiments were conducted on the same server, equipped with two AMD EPYC 7543 32-core CPUs and seven NVIDIA RTX A6000 GPUs. The deep learning-based model is trained using the GPUs, while the Algo model and our methods are trained using only the CPUs. The EmbedTextNet model takes approximately 30 minutes to train on relatively small GloVe word embeddings.
In contrast, the training time for our WordFS-P-wP and WordFS-S-wP models is less than 19 seconds and 13 seconds, respectively. Our methods without post-processing require even less time — the training of the WordFS-P-woP model is completed within one second. For more significant pre-trained word embeddings, it takes hours to train the EmbedTextNet model. However, our WordFS-P-wP and WordFS-S-wP methods only require 1-2 minutes, making them hundreds of times faster. Our WordFS-P-woP model, without post-processing, consistently completes in around one second, making it thousands or ten thousand times faster than the deep learning-based method. Besides, our method is significantly faster and outperforms UMAP. Because finding the K nearest neighbors may take a long time in a high-dimensional word vector space.
Additionally, our methods are more efficient than the PCA-based method because they employ PPA before and after the PCA, while we only apply it once before feature selection. The PPA consumes a substantial portion of the training time compared to the feature selection module. Our model, without the post-processing, can achieve even less time. The reason is the post-processing and PCA must be done on the entire vocabulary. However, our feature selection method only focuses on a subset of words that appear in the aggregated word similarity dataset, making the procedure very fast. And the module’s processing time will not be significantly affected by the vocabulary size.
In conclusion, Our methods take only a fraction of the time compared to deep learning methods and less than half the time compared to existing PCA-based methods. The experimental results demonstrate that our methods are much more efficient than existing methods.
@@ 5 Conclusion and Future Work
  • [13] Leland McInnes, John Healy, and James Melville. Umap: Uniform manifold approximation and projection for dimension reduction. arXiv preprint arXiv:1802.03426, 2018.